OpenStudy (anonymous):

integration X^3(x^2+4)^1/2 using subtitution or by part which on is u

OpenStudy (anonymous):

First you desegregate.

OpenStudy (anonymous):

Do a trig sub. \(x = 2\tan(\theta)\)

OpenStudy (anonymous):

what is desegregate???

OpenStudy (anonymous):

It's over your head... Do the substution \(x = 2\tan(\theta), dx=2\sec^2(\theta)d\theta \)

OpenStudy (anonymous):

@arifadli8 Do you know how to do this trig sub?

OpenStudy (anonymous):

i know a little about a trig sub

OpenStudy (anonymous):

\[ \int x^3\sqrt{x^2+4} dx = \int [2\tan(\theta)^2]^3\sqrt{[2\tan(\theta)]^2+4]}\sec^2(\theta)d\theta \]

OpenStudy (anonymous):

Use \(\tan^2(\theta)+1 = \sec^2(\theta)\) in the radical.

OpenStudy (anonymous):

Can you try it?

OpenStudy (anonymous):

i will.. but right know i want to understand about desegregate. if u have a link about desegregate. please give me

OpenStudy (anonymous):

I'll tell you about it once you finish integration, as a reward. Ok?

OpenStudy (anonymous):

why (2tan^2)^3... its should (2 tan)^3 @Wio please explain

OpenStudy (anonymous):

@Meepi can u explain

OpenStudy (anonymous):

You don't need to use the substitution \(x = \tan \theta \) for this one: \[\int x^3\sqrt{x^2 + 4} \, \mbox{d}x\] use the substitution \(u = x^2\), \(\frac{1}{2} \mbox{d}u = x\, \text{d}x\): \[\int u \sqrt{u + 4} \frac{1}{2}\, \text{d}u = \frac{1}{2} \int u \sqrt{u + 4}\, \text{d}u\] I'm pretty sure you can solve this one, just use the substitution v = u + 4 :) At the end substitute back u + 4 for v, then x^2 for u

OpenStudy (anonymous):

\(x = 2 \tan \theta\), not \(x = \tan \theta\), that was a typo :p But you don't need to use it anyway so if you don't know it yet don't worry about it

OpenStudy (anonymous):

You could probably do it in one substitution as well if you let u = x^2 + 4 um \[\int x^2 \sqrt{x^2 + 4} x dx\] u = x^2 + 4, du = 2x dx \[\frac{1}{2}\int (u - 4)\sqrt{u}du\] \[\frac{1}{2}\int u^{3/2} du - 2\int u^{1/2} du\]